OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-3, -1) with a slope of 4.

OpenStudy (istim):

You should post the equations.

OpenStudy (anonymous):

A) y - 1 = 4(x - 3) B) y + 1 = 4(x + 3) C) y - 3 = 4(x + 1) D) y + 3 = 4(x - 1)

OpenStudy (anonymous):

y+1=4(x-3) y=4x-4

OpenStudy (anonymous):

Thanks so much